# Connection between gravity and quantum mechanics: an algebraic approach to cosmological issues

### v1

@@ Abstract
This study presents an algebraic framework to explore the fundamental relationship between gravity and quantum mechanics, with particular emphasis on the role of primordial black holes (PBHs) in cosmology. Through the concept of self-gravitating condensed light, specifically in the form of a photon Bose-Einstein condensate, this work examines the quantum attributes of PBHs and their implications for early universe dynamics, baryogenesis, and the formation of large-scale structures. The model also investigates quantized characteristics of black holes, such as mass, entropy, and temperature, suggesting that quantum processes are fundamental to black hole mechanics. By addressing issues like the cosmological constant problem and the information loss paradox, this work provides insights into Planck-scale physics and proposes that PBHs may serve as a bridge between quantum theory and general relativity. This study ultimately posits that understanding black hole physics is essential to resolving major cosmological and astrophysical paradoxes for the ultimate unification of quantum mechanics with gravity.
@@ I Introduction
This paper aims to describe how primordial black holes are significant in the formation and evolution of the observable universe. Besides, it is trying to solve and interpret numerous questions and paradoxes relevant to Planck-scale physics, Big Bang cosmology, baryogenesis, the mechanism of Supermassive black holes (SMBHs), the early clumpiness of the universe, large-scale structures, such as supervoids and superclusters, and two-stage accelerated expansion; the horizon problem, cosmological constant problem, and the information-loss paradox that has arisen in observational cosmology and astrophysics. We shall demonstrate that the realistic physics of black holes is relevant to understanding fundamental cosmology problems. To clarify the above issues, let us first introduce the connection between the physics of primordial black holes and Bose-Einstein condensates at the Planck scale limit. This will be the main subject of the paper.
The fundamental discovery of a previously unknown state of matter, self-gravitating condensed light [1, 2], played an outstanding role in our research. We are discussing the quantized photon Bose-Einstein condensate [2, 3, 4], a phenomenon predicted approximately a century ago. Let us discuss the quantum nature of self-gravitating condensed light.
Within a vacuum, photons fail to be involved in direct interaction according to conventional quantum electrodynamics (QED). Photons often behave as though they interact via effective mass or higher-order quantum corrections (e.g., via virtual particle-antiparticle pairs in quantum loops). However, this may change in extreme situations like nonlinear media or close to strong gravitational fields. A self-gravitating system is possible because of these interactions, which allow photons to produce their gravitational fields potentially. The Euler-Heisenberg Lagrangian [1], which characterizes non-linear interactions at high energy scale QED, is the source of the idea of photon-photon scattering. Photons in such regimes may change the spacetime curvature around them and behave as gravity sources via self-interactions, which significantly change the behavior of light.
Bose-Einstein condensation (BEC), primarily associated with massive bosons such as atoms, can also be applicable to photons under specific conditions. Photons, as bosons, are capable of occupying identical quantum states, resulting in a macroscopic quantum phenomenon characterized by the presence of numerous photons in a coherent, condensed state. Considering the influence of gravity, these condensed photons may have the capacity to establish a self-gravitating system. Photons confined in optical cavities or thermalized through interactions with matter can attain BEC at temperatures above absolute zero [2]. At high energy densities, a self-bound photon configuration may theoretically emerge, similar to the formation of gravitationally bound systems such as stars or planets from massive particles.
The self-gravitating condensed light hypothesis introduces the idea of photon stars, which are gravitationally bound systems composed entirely of photons. Such theoretical objects are also known as geons, as they were conceived as solutions to Einstein’s field equations. Concentrated electromagnetic radiation would induce enough gravity to prevent the dispersion of photons [5]. Such stars are stable because the gravitational attraction is balanced by quantum pressure from photons in a Heisenberg uncertainty relation analogous to that for neutron stars, stabilizing the latter against collapse by neutron degeneracy pressure.
There is a balance between gravitational forces and quantum mechanical effects, especially the Heisenberg uncertainty principle, that makes a condensed light system stable in self-gravitating quantum frameworks. Quantum pressure, arising from the uncertainty in photon momentum and position, prevents gravitational collapse, similar to degeneracy pressure in white dwarfs and neutron stars. In semiclassical gravity [6], where quantum fields interact with classical spacetime, self-gravitating photon condensates may become stable in some situations. This gives us clues about where quantum field theory and general relativity meets. It might help us learn more about quantum gravity, cosmology, and high-energy astrophysics by showing us how photons behave in strong gravitational fields and how quantum effects happen in spacetime that is curved [7].
The early universe might have existed as a dense, “grainy” spacetime, possibly formed from a Planck-scale photon condensate, as proposed by quantum gravity theories [8, 9]. At this scale, spacetime is believed to be discrete rather than continuous, a notion supported by approaches like Loop Quantum Gravity(LQG) and string theory. These “grainy” structures could reflect quantum fluctuations, suggesting that photons formed a BEC under extreme energy conditions [10]. This condensate, where photons share a collective quantum state, may have acted as the foundational structure of spacetime. The Planck photon 3D condensate concept posits that at extremely high energies, especially near the Planck scale, photons may aggregate to form a BEC [8]. According to the holographic principle, which holds that black hole surfaces encode higher-dimensional information, PBHs can also be considered 2D photon condensates or ”membranes” that exhibit quantum characteristics [11].
The quantum characteristics of self-gravitating condensed light inherently lead to the quantization of the geometrical, gravitational, thermodynamical, informational, and radiational properties of black holes, which are described as follows:
The geometry of black holes is linked to the discrete characteristics of spacetime at the Planck scale. In quantum gravity theories such as LQG, spacetime is characterized as discrete, consisting of distinct “quanta” of space. This framework results in the inherent quantization of geometrical properties [12], including area and volume. In the context of black holes, this indicates that the surface area of the event horizon is quantized in units of the Planck area.
The gravitational field, as articulated by the Einstein field equations, is affected by quantum corrections at small scales [13]. In models of self-gravitating light, quantum fluctuations affect the gravitational field made by condensed photons. This creates a quantized spectrum of gravitational effects, especially in very dense objects like black holes.
Black hole thermodynamics, as articulated by Bekenstein-Hawking entropy [14, 15], demonstrates that black holes possess discrete energy levels, similar to those found in atomic systems. The entropy of a black hole, which is proportional to the area of its event horizon, is quantized in units of the Planck area. This is derived from the quantum mechanical interpretation of black hole horizons, in which microstates represent various quantum configurations of the horizon. Additionally, it should be noted that in his early works, S. Hawking et al. [16, 17] provided particular focus to the exceptional role of PBHs in his cosmological research.
The quantum features associated with black holes result in the quantization of information storage and radiation [18], especially as demonstrated by Hawking radiation. Black holes emit radiation in discrete quanta, and the information capacity of a black hole is proportional to the surface area of its event horizon. The quantization of information is directly connected to the black hole information paradox, wherein the loss of information through Hawking radiation presents an unresolved issue in quantum gravity. The quantized characteristics of black holes illustrate the profound relationship between quantum mechanics and gravity, particularly in extreme conditions where both theories are anticipated to merge into a unified framework.
On the other hand, V. Mukhanov and J. D. Bekenstein developed their research on quantized black holes, enhancing the comprehension of quantum phenomena in black hole physics [19, 20, 21]. Mukhanov’s research concentrated on quantum fluctuations during inflation [22], which subsequently proved essential to the hypothesis of cosmic structure development. His investigation on quantized black holes fostered discourse on the quantum characteristics of black hole horizons and the possibility of PBHs providing an understanding of quantum gravity. It should also be noted that recently, a growing number of astrophysicists contend that dark matter is comprised of microscopic PBHs with asteroid-like masses. A comprehensive review is available in Refs. [23, 24].
Recent remarkable discoveries, including Einstein’s gravitational waves [25, 26, 27, 28], “impossible” early galaxies and quasars [29, 30, 31, 32], and the 2D photon Bose-Einstein condensate with rest energy [2, 3, 35, 36], may offer a strong experimental and observational foundation for what John Wheeler famously described as the “fiery marriage of general relativity with quantum theory.” Let us discuss “impossible” early galaxies and quasars, and the 2D photon Bose-Einstein condensate with rest energy in brief.
Extremely large galaxies and quasars found in the early universe can pose serious questions to present-day models of cosmology. Observations of galaxies and quasars by advanced telescopes, like JWST, show that they have high masses, luminosities, and metallicities that go against current theories about how galaxies form and evolve. These theories are mostly about redshifts greater than $10$ [29, 30]. These so-called “impossible early galaxies” formed less than $500$ million years after Big Bang, indicating star formation rates and galaxy-assembly processes that exceed the predictions of conventional models. This indicates rapid star formation and galaxy assembly beyond the expectations of the standard Lambda Cold Dark Matter ($\Lambda$CDM) model, which actually predicts a gradual build-up of structure. Some galaxies appear massively, to the tune of $10^{9}$ solar masses within $300$ million years, and indicate accelerated star formation or unknown mechanisms [31]. Early quasars that are powered by SMBHs of mass $>10^{9}$ solar masses at redshift $>6$ also pose a challenge, since such black holes should take longer to form at the limits of accretion rates. Recent discoveries of quasars at redshifts $>7$ [32], challenge existing limits, and suggest the presence of “black hole seeds” that have significantly larger mass than those of consideration in earlier models or alternative growth mechanisms, including direct collapse scenarios. These results therefore demand reconsideration of $\Lambda$CDM or alternative theories, such as PBHs, non-standard dark matter interactions, or novel physics in the early universe.
Additionally, Belgiorno et al. [33] describe the analogous Hawking radiation that includes Bose-Einstein condensates and surface gravity waves. The examination of near-horizon phenomena in more controlled laboratory settings can be achieved through these systems that simulate analogous Hawking radiation. This facilitates investigations into grey-body factors, thermality, and dispersive effects, contributing to the experimental validation of Hawking radiation. The recent study conducted by Agullo et al. [34] examines the generation of entangled particle pairs within the context of 2D Bose-Einstein condensate analogs in expanding cosmological models. They have pointed out the significance of entanglement as a critical quantum indicator of pair creation and showcased experimental advancements in the potential to simulate quantum phenomena pertinent to cosmological contexts. These results converge to explain the influences of quantum gravity and its implications for the dynamics of the early universe and the formation of large-scale structures.
@@
On the other hand, the concept of a two-dimensional (2D) photon Bose-Einstein condensate introduces a novel approach to studying the quantum statistical behavior of light by combining quantum optics, condensed matter physics, and thermodynamics. Unlike traditional atomic BECs, where atoms condense at low temperatures, photons—typically massless—can achieve a condensed state by simulating an effective mass within a planar optical microcavity. This is achieved by confining photons between curved mirrors, creating a 2D environment where they behave similarly to massive particles. A heat reservoir, often via dye molecules, allows the photons to reach thermal equilibrium, forming a ”rest energy” and enabling condensation at reduced temperatures [2, 35]. As the system cools, photons transition to the lowest energy mode, forming a coherent quantum state [36]. This 2D photon BEC platform offers significant insights into fundamental physics, potentially impacting low-threshold laser technology, quantum information, and the study of thermodynamic properties in non-equilibrium quantum systems [35]. Recently, Mazur and Mottola [37] introduced the concept of Gravitational Condensate Stars as a stable alternative to black holes. In their work, they present a compact, non-singular solution to Einstein’s equations, suggesting it as a possible endpoint for gravitational collapse. Unlike black holes, this type of collapsed star aligns with quantum theory, remains thermodynamically stable, and avoids the information paradox.
Integrating general relativity and quantum mechanics is crucial to deepening our understanding of the universe and enhancing space exploration. While effective in their domains, these theories must unify to fully explain extreme cosmic phenomena like black holes and the early universe. Gravitational waves (spacetime ripples) and quantum probability waves are fundamentally different yet share a geometric link, suggesting a pathway for unification. This approach involves analyzing phenomena, especially black holes, through wavelengths and geodesics, bridging the two frameworks. A key objective is to identify a quantum “background substance” that balances black holes’ intense gravitational forces with quantum repulsion, shedding light on effects like Hawking radiation and quantum tunneling. This could ultimately deepen our grasp of the universe across both large and small scales.
@@
This manuscript is structured in the following manner: In section II, we derived the algebra (namely Einstein’s algebra) utilizing the membrane formalism, establishing a correspondence between the primordial black holes and the two-dimensional condensate photons. In section III, we examine the essential foundational principles of physics at the Planck scale. In section IV, we elucidate the interconnections between energy, temperature, and entropy of black holes when they approach the Planck scale limit. The discussion of primordial black holes’ lifetime and luminosity, in conjunction with the information loss paradox, is presented in Section V. Subsequently, we outlined the process of free light accretion via quantum absorption, as well as the origins of disks and jets in section VI. In section VII, we analyze the conservation law of quantum information and the energies associated with gravitational radiation from colliding and merging black holes, which serve as a driving force behind the accelerated expansion of the universe. In section VIII, we examine the intricate details of Planck-scale phenomena, the dynamics of the hot Big Bang, and the initial phase of the universe’s accelerated expansion. In section IX, we derived the connection between the Planckian Era and the Baryogenesis Epoch, addressing the issue of supermassive black hole formation and the primordial clumping of the universe. In section X, we analyze baryogenesis via the mechanism of primordial black hole production. In section XI, we established the methodology for the formation of contemporary dark matter and the large-scale structure of the universe. The final section, XII, serves as the discussion and conclusion of this study.
@@ II The sketches of “Einstein’s algebra”
In this section, we begin with Albert Einstein’s final work in 1956 [38], where he articulated a concluding statement that reflects his core methodological principle: “… a finite system of finite energy can be completely described by a finite set of quantum numbers. This does not seem to align with a continuum theory and must lead to an attempt to find a purely algebraic theory for describing reality. However, nobody knows how to establish the foundation for such a theory.”
@@
The previously mentioned concepts may also apply to 2D photon Bose-Einstein condensates (BEC) with rest energy in various phenomena [2, 3, 45, 46]. The idea of a 2D photon BEC with rest energy can be linked to the early universe, particularly concerning PBHs, analog gravity, and related topics [47, 48, 49, 50]. This correlation can be framed within a physical model where PBHs are represented as 2D spherical photon condensates contained within their own gravitational fields. Such a model naturally leads to a geometrical equation that directly connects the Compton wavelength, $\lambda_{sm}$, of condensed light (quantum theory) with the geodesic length (general relativity) [4]:
@@
where $R_{s}$ is Schwarzschild radius. Based on the above discussions, we can intuitively say that spacetime instructs condensed light quanta on how to curve, while condensed light quanta instructs spacetime on how to quantize. According to the definition of the Compton wavelength and the expression for Schwarzschild radius, we may express it as
$\lambda_{sm}=$ $\frac{2\pi\hbar}{M_{sm}c}$
$2\pi R_{s}=$ $\frac{4\pi GM_{s}}{c^{2}}$ (2)
Using the above relationship Eqs. (1) and (2), we promptly determine the mass and rest energy of each condensed photon:
@@
along with their aggregate quantity:
@@
expressed in terms of the Planck mass $M_{p}\leavevmode\nobreak\ (=\sqrt{\frac{\hbar c}{G}})$, Planck energy $E_{p}$, and Planck length $l_{p}\leavevmode\nobreak\ (=\sqrt{\frac{\hbar G}{c^{3}}})$. The fundamental and continuous relationships may enable us to explore the unknown facets of Planck-scale physics. Considering that the quantity $N_{s}$ of condensed photons cannot be less than two, we readily derive a significant expression for $N_{s}$, utilizing the natural quantum number $n_{s}$ as
@@
Relying solely on established principles of physics, one can readily determine the quantized characteristics of PBHs - such as mass $M_{s}$, rest energy $E_{s}$, size $R_{s}$, entropy $S_{s}$, temperature $T_{s}$, quantum information quantity $I_{s}$, longevity $\tau_{s}$, luminosity $L_{s}$, among others—as well as the fundamental principles governing their formation, evolution, and eventual decay. It is important to note that all these quantized features of PBHs depend exclusively on Planck units and are influenced by various powers of natural quantum numbers [4]. From the previously established quantum relations we get,
@@
$I_{s}=N_{s}=2n_{s}(qubit);\leavevmode\nobreak\ \tau_{s}=\frac{4\pi t_{p}n_{s}^{\frac{3}{2}}}{3};\leavevmode\nobreak\ L_{s}=\frac{L_{p}}{4\pi n_{s}},$ (6)
@@
In this discussion, we focus on the quantum mechanical principles underlying the phenomenon of a black hole (PBHs). By rephrasing $n_{s}$ in Planck scale units, we can enable a direct comparison of our results with the findings drawn by Bekenstein [51] and Hawking [14]:
@@
where we define $S_{B}\leavevmode\nobreak\ (=\frac{k_{B}c^{3}ln2}{8\pi G\hbar}A)$ to represent Bekenstein entropy [51], $S_{H}\leavevmode\nobreak\ (=\frac{k_{B}c^{3}}{4G\hbar}A)$ to signify Hawking entropy [14], and $T_{H}\leavevmode\nobreak\ (=\frac{\hbar c^{3}}{8\pi GMk_{B}})$ to denote Hawking temperature.
Our findings show that our outcomes closely match those that Bekenstein and Hawking predicted, indicating that we are moving in the right direction. However, our findings were derived not from intuitive assumptions, but through clear and exact physical principles, relationships, and an adequate experimental basis.

### v2

# Connecting Gravity and Quantum Physics: Primordial Black Holes and the Evolution of the Universe

@@ Abstract
This study presents a novel framework to explore the fundamental relationship between gravity and quantum mechanics, with particular emphasis on the role of primordial black holes (PBHs) in cosmology. Through the concept of self-gravitating condensed light, specifically in the form of the experimentally discovered photon Bose-Einstein condensate, this work examines the quantum attributes of PBHs and their implications for early universe dynamics, baryogenesis, and the very early formation of galaxies and large-scale structures, established by JWST data. The model also investigates quantized characteristics of PBHs, such as mass, entropy, and temperature, suggesting that quantum processes are fundamental to black hole mechanics. By addressing issues like the cosmological constant problem and the information loss paradox, dark matter, and dark energy, this work provides insights into Planck-scale physics and proposes that PBHs may serve as a bridge between quantum theory and general relativity. This study ultimately posits that understanding PBH physics is essential to resolving major cosmological and astrophysical paradoxes for the ultimate unification of quantum mechanics with gravity.
@@ I Introduction
Gravity and quantum mechanics are widely regarded as the two pillars of modern physics, each describing the universe with unparalleled success but in fundamentally different realms. Gravity, as encapsulated by Einstein’s theory of General Relativity, governs the macroscopic world: the motion of planets, the dynamics of galaxies, and the curvature of spacetime itself. Its elegance lies in the geometric framework, where massive objects dictate the shape of spacetime, and spacetime, in turn, guides the motion of these objects.
Quantum mechanics, on the other hand, is the language of the microscopic world. It describes the behavior of particles at the atomic and subatomic scales governed by probabilities, wave functions, and inherent uncertainty. This framework has given rise to astonishing technologies and profound insights into the nature of matter and energy, yet it operates in a domain where classical notions of reality are often defied.
Despite their individual successes, gravity and quantum mechanics resist unification. General relativity treats spacetime as a smooth continuum, while quantum mechanics thrives on discreteness and probabilistic events. Attempts to bridge these paradigms often encounter insurmountable challenges: mathematical inconsistencies, infinite quantities, and a lack of experimental evidence at the scales where their interplay would manifest.
This profound incompatibility underscores the uniqueness of both frameworks and raises one of the most profound questions in physics: How can we reconcile the smoothness of spacetime with the granularity of quantum mechanics? The pursuit of a quantum theory of gravity, whether through string theory, loop quantum gravity, or other approaches, continues to push the boundaries of our understanding. It is not merely a quest for unification, but a journey to uncover a deeper, more fundamental description of reality.
The need for a quantum theory of gravity arises from the limitations of our current theoretical frameworks when applied to extreme conditions in the universe. General Relativity successfully describes the behavior of spacetime and gravity on large scales, but it breaks down under extreme conditions, such as within black holes or at the very beginning of the universe during the Big Bang. At these scales, quantum effects become significant, and a theory that combines both quantum mechanics and gravity is essential to provide a complete description of nature.
Without a quantum theory of gravity, our understanding of phenomena like the singularities in black holes or the quantum fluctuations that seeded the structure of the universe remains incomplete. Additionally, the unification of gravity with the other fundamental forces—electromagnetic, weak, and strong—requires a consistent framework that integrates all forces under the same quantum principles. A quantum theory of gravity would also have profound implications for the nature of spacetime, that helps revealing its granular structure and resolving paradoxes like the information loss problem in black holes.
The development of a quantum gravity theory is not only a theoretical pursuit but also a step toward addressing practical questions in cosmology, high-energy physics, and even quantum computing. It promises to bridge the gap between the macroscopic and microscopic and provides a unified view of the universe from the smallest particles to the vast cosmos.
In this regard, we want to mention that, the study of primordial black holes (PBHs) and two-dimensional (2D) photon Bose-Einstein condensate (BEC) are crucial for connecting gravity with quantum mechanics, as they naturally combine extreme gravitational and quantum mechanical effects. Formed shortly after the Big Bang, PBHs probe the interplay of quantum fluctuations and gravitational collapse under early-universe conditions, which offers unique insights into Planck-scale physics. Their potential evaporation via Hawking radiation provides a direct test of quantum field theory in curved spacetime, while their small masses allow us to validate the quantum gravity phenomena like black hole thermodynamics, information loss paradoxes, and potential remnants. PBHs also constrain cosmological models, link to dark matter hypotheses, and contribute to gravitational wave generation [1, 2]. By studying PBHs, we can investigate quantum gravity theories, explore the statistical nature of black hole entropy, and seek observable signatures like gamma-ray bursts or microlensing, making them a potential candidate that may help us to unify quantum mechanics and general relativity[3].
Along with this, the recent developments in PBHs research have shown its substantial cosmic consequences, as emphasized by Carr et al. [4, 5]. Their research offers convincing indirect evidence that PBHs, especially those with extensive mass distribution, could illustrate several astrophysical anomalies, such as the emergence of ultra-faint dwarf galaxies, the dynamical heating of the Galactic disk, and high-redshift quasars, presenting alternatives to the conventional cold dark matter paradigm. This corresponds with the increasing recognition of PBHs as significant factors in early cosmology, perhaps acting as progenitors for supermassive black holes (SMBHs) and facilitating the construction of cosmic structures. The authors of [4] present a “positivist perspective,” highlighting the unifying potential of PBHs in resolving outstanding astrophysical queries, further substantiated by new associations with photon Bose-Einstein condensates. Comprehensive research on PBHs and their cosmic significance, including their birth and growth, is documented in the following Refs. [6, 7, 8, 9, 10]
On the other hand, BEC comprising a two-dimensional condensate photon gas [11, 12] offers a distinctive platform for investigating and evaluating ideas in quantum gravity, functioning as a quantum-coherent system that reflects characteristics of spacetime and gravitational processes. In such a condensate, the alteration of the effective potential encountered by photons can replicate curved spacetime, generating analogs of event horizons and gravity wells[13]. This facilitates the examination of phenomena such as Hawking radiation inside a regulated laboratory setting [14] under analog gravity, providing experimental insights into quantum field theory in curved spacetime[15].
The macroscopic coherence of the photon BEC also embodies characteristics fundamental to quantum gravity, facilitating the investigation of how quantum phenomena may influence the formation of spacetime and gravity [15]. Quantum fluctuations, especially evident in two-dimensional systems, offer additional avenues for exploring dynamics similar to those of the early universe[16, 17]. The thermal and entropic characteristics of the condensate align with the laws of black hole thermodynamics, providing an analogy to the Bekenstein-Hawking entropy [18, 19] and enhancing our comprehension of quantum geometry.
Through the analysis of these systems, we can investigate the interaction between quantum mechanics and gravity, simulate emergent spacetime, and evaluate notions such as coherence, thermodynamics, and quantum field behavior. The 2D photon BEC connects theoretical physics with experimental practices, and may offer a profound opportunity to illuminate the path towards establishing a connection between quantum mechanics and general relativity.
@@
Thus, the concept of a 2D photon BEC introduces a new approach to study the quantum statistical behavior of light by combining quantum optics, condensed matter physics, and thermodynamics. Unlike traditional atomic BECs, where atoms condense at low temperatures, photons—typically massless—can achieve a condensed state by simulating an effective mass within a planar optical microcavity. This is achieved by confining photons between curved mirrors, creating a 2D environment where they behave similarly to massive particles. A heat reservoir, often via dye molecules, allows the photons to reach thermal equilibrium, forming a ”rest energy” and enabling condensation at reduced temperatures [13, 20]. As the system cools, photons transition to the lowest energy mode, forming a coherent quantum state [21]. This 2D photon BEC platform offers significant insights into fundamental physics, potentially impacting low-threshold laser technology, quantum information, and the study of thermodynamic properties in non-equilibrium quantum systems [20]. Recently, Mazur and Mottola [22] introduced the concept of Gravitational Condensate Stars as a stable alternative to black holes. In their work, they present a compact, non-singular solution to Einstein’s equations, suggesting it as a possible endpoint for gravitational collapse. Unlike black holes, this type of collapsed star aligns with quantum theory, remains thermodynamically stable, and avoids the information loss paradox.
Besides this, Recent remarkable discoveries, including Einstein’s gravitational waves [23, 24, 25, 26], “impossible” early galaxies and quasars [27, 28, 29, 30], and the 2D photon Bose-Einstein condensate with rest energy [11, 13, 20, 21], may offer a strong experimental and observational foundation for the existence of a quantum gravity theory , which John Wheeler famously described as the “fiery marriage of general relativity with quantum theory.”
In this study, the objective is to integrate the aforementioned concepts to establish a connection between the geometric nature of gravity and the discrete properties of quantum mechanics through an algebraic framework, utilizing the unique perspectives offered by PBHs and the two-dimensional self-gravitating photon BEC. This approach is unique in its synthesis of astrophysical phenomena with aid from quantum statistical mechanics, providing a fertile ground to investigate the quantum aspects of gravity.
Our study aims to illustrate the significance of PBHs in the formation and evolution of the observable universe. It may act as the key factor to solve major questions and paradoxes relevant to Planck-scale physics, such as Big Bang cosmology, the physical background of dark matter and dark energy, baryogenesis, the formation mechanism of SMBHs, early clumpiness of the universe, large-scale structures, such as super-voids and superclusters, and two-stage accelerated expansion of the Universe; the horizon problem, cosmological constant problem, and the information-loss paradox that has arisen in observational cosmology and astrophysics.
By analyzing these specific problems we want to develop a common understanding of how algebraic techniques might shed light on the fundamental properties of spacetime and its quantum features. This method allows us to address theoretical issues and set a framework for future experimental validation.
@@
This manuscript is structured as follows: In section II, we derived the algebra (namely Einstein’s algebra) utilizing the membrane formalism, establishing a correspondence between the PBHs and the two-dimensional condensate photons. In section III, we examine the essential foundational principles of physics at the Planck scale. Starting from Section IV, we will utilize our newly developed formalism to explore a diverse array of phenomena, including black hole thermodynamics, luminosity, lifetime, the hot Big Bang, accelerated expansion, and baryogenesis, testing the success of our method to solve some previously unsolved problems of cosmology. In Section IV we shall elucidate the interconnections between energy, temperature, and entropy of black holes when they approach the Planck scale limit. The discussion of PBHs’ lifetime and luminosity, in conjunction with the information loss paradox, is presented in Section V. Subsequently, we outlined the process of free light accretion via quantum absorption, as well as the origins of disks and jets in Section VI. In section VII, we analyze the conservation law of quantum information and the energies associated with gravitational radiation from colliding and merging of black holes, which serve as a driving force behind the accelerated expansion of the universe. In section VIII, we examine the intricate details of Planck-scale phenomena, the dynamics of the hot Big Bang, and the initial phase of the universe’s accelerated expansion. In section IX, we derived the connection between the Planckian Era and the Baryogenesis Epoch, addressing the issue of supermassive black hole formation and the primordial clumping of the universe. In section X, we analyze baryogenesis via the mechanism of primordial black hole production. In section XI, we established the methodology for the formation of contemporary dark matter and the large-scale structure of the universe. The final section, XII, serves as the conclusion of this study.
@@ II The sketches of “Einstein’s algebra”
We begin this section with Albert Einstein’s final work in 1956 [31], where he articulated a concluding statement that reflects his core methodological principle: “… a finite system of finite energy can be completely described by a finite set of quantum numbers.” This does not seem to align with a continuum theory and must lead to an attempt to find a purely algebraic theory for describing reality. However, nobody knows how to establish the foundation for such a theory.
Inspired by Einstein’s vision, we aim to develop a theory rooted purely in algebra and try to bridge the seemingly insurmountable gap between gravity and quantum mechanics.
@@
The above-mentioned concepts may also apply to 2D photon Bose-Einstein condensates (BEC) with rest energy (arise due to condensation) in various phenomena [11, 13, 38, 39]. The idea of a 2D photon BEC with rest energy can be linked to the early universe, particularly concerning PBHs, analog gravity, and related topics [40, 41, 42, 43]. This correlation can be framed within a physical model where PBHs are represented as 2D spherical photon condensates contained within their own gravitational fields [44]. Such a model naturally leads to a geometrical equation that directly connects the Compton wavelength, $\lambda_{sm}$, of condensed light (quantum theory) with the geodesic length (general relativity) [44]:
@@
where $R_{s}$ is Schwarzschild radius. From the preceding considerations, we can intuitively say that spacetime instructs condensed light quanta on how to curve, while condensed light quanta instructs spacetime on how to quantize. The expression for Schwarzschild radius in term of mass ($M_{s}$) of the PBH, may be expressed as
$R_{s}=\frac{2GM_{s}}{c^{2}}$ (2)
In a 2D photon BEC, photons can exhibit behavior analogous to massive particles due to the effective mass they acquire within the system, even though photons are intrinsically massless in free space. This effective mass arises from the interaction of photons with the medium and the trapping potential used to create the condensate. Relating this effective mass($M_{sm}$) to its Compton wavelength ($\lambda_{sm}$) we can write,
$\lambda_{sm}=$ $\frac{2\pi\hbar}{M_{sm}c}$ (3)
Using the above relationship Eqs. (1),(2) and (3), we promptly determine the effective mass and rest energy of each condensed photon in terms of Compton wavelength ($\lambda_{sm}$) and Schwarzschild radius($R_{s}$):
@@
expressing the above quantities in terms of condensate photon number ($N_{s}$):
@@
expressed in terms of the Planck mass $M_{p}\leavevmode\nobreak\ (=\sqrt{\frac{\hbar c}{G}})$, Planck energy $E_{p}$, and Planck length $l_{p}\leavevmode\nobreak\ (=\sqrt{\frac{\hbar G}{c^{3}}})$. In the above Eq.(5), $M_{s}$ and $E_{s}$ stands for total rest mass and the rest energy of the condensate photon respectively. These fundamental and consistent relationships may enable us to explore the unknown facets of Planck-scale physics. Considering that the quantity $N_{s}$ of condensed photons cannot be less than two to form a bound state, we can readily derive a significant expression for $N_{s}$, utilizing the natural quantum number $n_{s}$ as
@@
Relying solely on established principles of physics, one can readily determine the quantized characteristics of PBHs - such as mass $M_{s}$, rest energy $E_{s}$, size $R_{s}$, area $A_{s}$, entropy $S_{s}$, temperature $T_{s}$, quantum information quantity $I_{s}$, longevity $\tau_{s}$, luminosity $L_{s}$, among others—as well as the fundamental principles governing their formation, evolution, and eventual decay. It is important to note that all these quantized features of PBHs depend exclusively on Planck units and are influenced by various powers of natural quantum numbers [44]. From the previously established quantum relations we get,
@@
$I_{s}=N_{s}=2n_{s}(bit);\leavevmode\nobreak\ \tau_{s}=\frac{4\pi t_{p}n_{s}^{\frac{3}{2}}}{3};\leavevmode\nobreak\ L_{s}=\frac{L_{p}}{4\pi n_{s}},$ (7)
@@
To facilitate a direct comparison with the foundational findings of Bekenstein [18] and Hawking [19] for the black hole entropy, we express $n_{s}$ in terms of the area $A_{s}$ of the black hole using Eq. (7). By substituting the expression for the Planck length $l_{p}$ into the formulation, we obtain a modified expression that illustrates the connection between quantum states and the geometric characteristics of the black hole as:
@@
where we define $S_{B}\leavevmode\nobreak\ (=\frac{k_{B}c^{3}ln2}{8\pi G\hbar}A)$ to represent Bekenstein entropy [18], $S_{H}\leavevmode\nobreak\ (=\frac{k_{B}c^{3}}{4G\hbar}A)$ to signify Hawking entropy [19], and $T_{H}\leavevmode\nobreak\ (=\frac{\hbar c^{3}}{8\pi GMk_{B}})$ to denote Hawking temperature. Note that historically Jacob Bekenstein was the first physicist who combined all four universal constants, $G,c,\hbar,k_{B}$ related to quantum mechanics and gravity in one consistent formula for the black hole entropy.
Looking at the result (8) we can state that our findings closely matched to those that Bekenstein and Hawking predicted, which indicates that we are moving in the right direction. However, our findings were derived not from intuitive assumptions, but through clear and exact physical principles of physics.